# A Survey on Knowledge Distillation of Large Language Models

### 20 Feb 2024 v1

@@ Abstract
This survey presents an in-depth exploration of knowledge distillation (KD) techniques within the realm of Large Language Models (LLMs), spotlighting the pivotal role of KD in transferring sophisticated capabilities from proprietary giants such as GPT-4 to accessible, open-source models like LLaMA and Mistral. Amidst the evolving AI landscape, this work elucidates the critical disparities between proprietary and open-source LLMs, demonstrating how KD serves as an essential conduit for imbuing the latter with the former’s advanced functionalities and nuanced understandings. Our survey is meticulously structured around three foundational pillars: algorithm, skill, and verticalization – providing a comprehensive examination of KD mechanisms, the enhancement of specific cognitive abilities, and their practical implications across diverse fields. Crucially, the survey navigates the intricate interplay between data augmentation (DA) and KD, illustrating how DA emerges as a powerful paradigm within the KD framework to bolster LLMs’ performance. By leveraging DA to generate context-rich, skill-specific training data, KD transcends traditional boundaries, enabling open-source models to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts. This work aims to provide an insightful guide for researchers and practitioners, offering a detailed overview of current methodologies in knowledge distillation and proposing future research directions. By bridging the gap between proprietary and open-source LLMs, this survey underscores the potential for more accessible, efficient, and sustainable AI solutions, fostering a more inclusive and equitable landscape in AI advancements. An associated Github repository is available at https://github.com/Tebmer/Awesome-Knowledge-Distillation-of-LLMs.
@@ 1 Introduction
Recognizing the disparities between proprietary and open-source LLMs, the field of NLP has witnessed a surge in the adoption of knowledge distillation techniques Gou et al. (2021); Gupta and Agrawal (2022) as a means to bridge this performance gap. Knowledge distillation, in this context, involves leveraging the more advanced, nuanced capabilities of larger, proprietary models like GPT-4 or Gemini as a guiding framework to enhance the competencies of open-source LLMs. This process is akin to transferring the ‘knowledge’ of a highly skilled teacher to a student, wherein the student (e.g., open-source LLM) learns to mimic the performance characteristics of the teacher (e.g., proprietary LLM). Compared to traditional knowledge distillation algorithms Gou et al. (2021), data augmentation (DA) Feng et al. (2021) has emerged as a prevalent paradigm to achieve knowledge distillation of LLMs, where a small seed of knowledge is used to prompt the LLM to generate more data with respect to a specific skill or domain Taori et al. (2023).
@@
A key aspect of this knowledge transfer is the enhancement of skills such as advanced context following (e.g., in-context learning Huang et al. (2022a) and instruction following Taori et al. (2023)), improved alignment with user intents (e.g., human values/principles Cui et al. (2023a), and thinking patterns like chain-of-thought (CoT) Mukherjee et al. (2023)), and deeper language understanding (e.g., machine reasoning Hsieh et al. (2023), semantic understanding Ding et al. (2023a), and code generation Chaudhary (2023)). These skills are crucial for the wide array of applications that LLMs are expected to perform, ranging from casual conversations to complex problem-solving in specialized domains. For instance, in vertical domains like healthcare Wang et al. (2023a), law LAW (2023), or science Zhang et al. (2024), where accuracy and context-specific knowledge are paramount, knowledge distillation allows open-source models to significantly improve their performance by learning from the proprietary models that have been extensively trained and fine-tuned in these areas.
@@
The benefits of knowledge distillation in the era of LLMs are multifaceted and transformative Gu et al. (2024). Through a suite of distillation techniques, the gap between proprietary and open-source models is significantly narrowed Chiang et al. (2023); Xu et al. (2023a) and even filled Zhao et al. (2023a), enabling the latter to achieve higher levels of performance and efficiency previously exclusive to their proprietary counterparts. This process not only streamlines computational requirements but also enhances the environmental sustainability of AI operations, as open-source models become more proficient with lesser computational overhead. Furthermore, knowledge distillation fosters a more inclusive and equitable AI landscape, where smaller entities and individual researchers gain access to state-of-the-art capabilities, encouraging wider participation and diversity in AI advancements. This democratization of technology leads to more robust, versatile, and accessible AI solutions, catalyzing innovation and growth across various industries and research domains.
@@ 2.1 Comparing Traditional Recipe
2.2 Relation to Data Augmentation (DA)
2.3 Survey Scope
Skill Distillation. This facet examines the specific competencies and capabilities enhanced through KD. It encompasses detailed discussions on context following Taori et al. (2023); Luo et al. (2023a), with subtopics like instruction following and retrieval-augmented generation (RAG) Capability. In the realm of alignment Mitra et al. (2023); Tunstall et al. (2023), the survey investigates thinking patterns, persona/preference modeling, and value alignment. The ‘agent’ category delves into skills such as Tool Using and Planning. NLP task specialization Dai et al. (2023a); Jung et al. (2023); Chaudhary (2023) is scrutinized through lenses like natural language understanding (NLU), natural language generation (NLG), information retrieval, recommendation systems, text generation evaluation, and code generation. Finally, the survey addresses multi-modality Liu et al. (2023b); Zhao et al. (2023b), exploring how KD enhances LLMs’ ability to interpret and integrate multiple forms of input, enriching their utility and applicability across various contexts.
@@ 2.3 Survey Scope
Self-Align Sun et al. (2024b), WizardLM Xu et al. (2023a), WizardCoder Luo et al. (2023b),
WizardMath Luo et al. (2023c), AugGPT Dai et al. (2023a), TDG He et al. (2023b), leaf, text width=31em ] ] [ Curation [ UltraChat Ding et al. (2023b), Phi-1 Gunasekar et al. (2023), Phi-1.5 Li et al. (2023a),
@@
GKD Agarwal et al. (2024), QuantGPT Tao et al. (2022a), LLM-QAT Liu et al. (2023a), , leaf, text width=31em ] ] [ Feedback [ CAI Bai et al. (2022a), WizardMath Luo et al. (2023c), UltraFeedback Cui et al. (2023a),
@@
TED Liang et al. (2023a), GKD Agarwal et al. (2024),BabyLlamaTimiryasov and Tastet (2023), leaf, text width=31em ] ] [ Reinforcement Learning [ CAI Bai et al. (2022a), UltraFeedback Cui et al. (2023a), WizardMath Luo et al. (2023c),
@@
WizardMath Luo et al. (2023c), Llama-GPT4 Peng et al. (2023a), , leaf, text width=31em ] ] [ Multi-turn Dialogue [ Vicuna Chiang et al. (2023), Baize Xu et al. (2023b), UltraLLaMA Ding et al. (2023b),
CAMEL Li et al. (2023b), OpenChat Wang et al. (2023c), Zephyr Tunstall et al. (2023), , leaf, text width=31em ] ] [ RAG Capbility [ KARD Kang et al. (2023a), SAIL Luo et al. (2023a), Self-RAG Asai et al. (2023), , leaf, text width=31em ] ] ] [ Alignment [ Thinking Pattern [ Selfee Ye et al. (2023), Orca Mukherjee et al. (2023), Orca 2 Mitra et al. (2023),
@@
Zephy Tunstall et al. (2023), UltraFeedback Cui et al. (2023a), , leaf, text width=31em ] ] [ Value [ CAI Bai et al. (2022a), Align Honesty Yang et al. (2023a), SANDBOX Liu et al. (2023c),
@@
RankZephyr Pradeep et al. (2023b), ExaRanker Ferraretto et al. (2023), , leaf, text width=31em ] ] [ Recommendation [ NDR Mysore et al. (2023), InstrcutRec Zhang et al. (2023b), ONCE Liu et al. (2023d), , leaf, text width=31em ] ] [ Text Generation Evaluation [ PandaLM Wang et al. (2023b), Prometheus Kim et al. (2024), InstructScore Xu et al. (2023d),
@@
Phi-1 Gunasekar et al. (2023), PERsD Chen et al. (2023a), MFTCoder Liu et al. (2023e),
WaveCoder Yu et al. (2024), Code Clean Jain et al. (2023), , leaf, text width=31em ] ] ] [ Multi-Modality [ LLaVA Liu et al. (2023b), SVIT Zhao et al. (2023b), LVIS-Instruct4V Wang et al. (2023e), Shikra Chen et al. (2023c),
@@ 3.1.2 Expansion
While the labeling approach is simple and effective, it faces certain limitations. Primarily, it is constrained by the scale and variety of the input data. In real-world applications, especially those involving user conversations, there are also concerns regarding the privacy of the data involved. To address these limitations, various expansion methods have been proposed Wang et al. (2022a); Taori et al. (2023); Chaudhary (2023); Si et al. (2023); Ji et al. (2023a); Luo et al. (2023c, b); Wu et al. (2023c); Sun et al. (2024b); Xu et al. (2023a); Guo et al. (2023c); Rozière et al. (2023); West et al. (2022). These methods take the demonstrations as seed knowledge and aim to expand a large scale and various data by in-context learning.
@@
Expansion techniques have been widely utilized to extract extensive instruction-following knowledge from teacher LLMs. Wang et al. (2022a) first introduces an iterative bootstrapping method, Self-Instruct, to utilize LLMs to generate a wide array of instructions based on several demonstrations sampled from 175 manually-written instructions. The newly generated instructions are then added back to the initial pool, benefiting subsequent expansion iterations. Subsequently, Taori et al. (2023) applies this expansion method to a more powerful teacher LLM, text-davinci-003, to distill 52K high-quality data. To improve the diversity and coverage during expansion, Wu et al. (2023c) and Sun et al. (2024b) prompt the teacher LLM to generate instructions corresponding to some specific topics. Xu et al. (2023a) propose an Evol-Instruct method to expand the instructions from two dimensions: difficulty (e.g. rewriting the question to be more complex) and diversity (e.g. generating more long-tailed instructions). This Evol-Instruct method is domain-agnostic and has been used to expand the distillation of coding Luo et al. (2023b) and math Luo et al. (2023c). Additionally, expansion methods can significantly augment NLP task datasets with similar samples, thereby enhancing task performance. For instance, AugGPT Dai et al. (2023a) leverages a teacher LLM to rephrase each sentence in the training samples into multiple conceptually similar, but semantically varied, samples to improve classification performance. Similarly, TDG He et al. (2023b) proposes the Targeted Data Generation (TDG) framework, which automatically identifies challenging subgroups within data and generates new samples for these subgroups using LLMs through in-context learning.
@@ 3.1.3 Data Curation
The pursuit of high-quality and scalable data generation in knowledge distillation from LLMs has led to the emergence of the Data Curation approach. This method arises in response to the limitations observed in both the Labeling and Expansion approaches. These methods often yield data of variable quality and face constraints in quantity. In Labeling, the seed knowledge is sourced from task datasets, leading to potential noise and dirty data. Meanwhile, in Expansion, the input $x$ is derived from seed demonstrations, which can result in homogeneous data when generated in large quantities. To overcome these challenges, the Data Curation method curates high-quality or large-scale data by extensive meta-information as seed knowledge Ding et al. (2023b); Gunasekar et al. (2023); Li et al. (2023a); Mar (2023); Liu et al. (2023e); Wei et al. (2023); Yu et al. (2024); Ye et al. (2022); Gao et al. (2023a); Yang and Nicolai (2023).
@@
Different studies primarily vary in their source and method of leveraging meta-information. UltraChat Ding et al. (2023b) effectively demonstrates the process of curating both high-quality and diverse data by distilled knowledge. They collect extensive meta-information across three domains: Questions about the World, Creation and Generation, and Assistance on Existing Materials. For example, under Questions about the World, they explore 30 meta-topics like ”Technology” and ”Food and Drink.” the teacher LLMs then use this meta-information to distill a broad array of instructions and conversations, achieving a substantial scale of 1.5 million instances. UltraChat stands out with its lexical and topical diversity. The UltraLLaMA model, fine-tuned on this data, consistently surpasses other open-source models. Another notable series, phi Gunasekar et al. (2023); Li et al. (2023a); Mar (2023), focuses on distilling smaller, high-quality datasets akin to ”textbooks.” Phi-1Gunasekar et al. (2023) experiments with synthesizing ”textbook quality” data in the coding domain. Their approach involves distilling clear, self-contained, instructive, and balanced content from LLMs, guided by random topics or function names to enhance diversity. The distilled data is a synthesis of 1 billion tokens of Python textbooks, complete with natural language explanations and code snippets, as well as 180 million tokens of Python exercises with solutions. Remarkably, the phi-1 model, despite its smaller size, outperforms nearly all open-source models on coding benchmarks like HumanEval and MBPP while being 10 times smaller in model size and 100 times smaller in dataset size. MFTCoder Liu et al. (2023e) utilizes hundreds of Python knowledge points as meta-information to create a CodeExercise Dataset. In contrast, Magicoder Wei et al. (2023) and WaveCoder Yu et al. (2024) get raw code collections from open-source code datasets, using this as meta-information for generating instructional data. In the context of NLU tasks, certain studies Ye et al. (2022); Gao et al. (2023a); Wang et al. (2021a) explore the use of labels as meta-information to synthesize corresponding samples for data augmentation. Similarly, in information retrieval tasks, there are efforts to utilize documents as meta-information for generating potential queries, thereby constructing large-scale retrieval pairs Bonifacio et al. (2022); Meng et al. (2023).

### 23 Feb 2024 v2

@@ Abstract
In the era of Large Language Models (LLMs), Knowledge Distillation (KD) emerges as a pivotal methodology for transferring advanced capabilities from leading proprietary LLMs, such as GPT-4, to their open-source counterparts like LLaMA and Mistral. Additionally, as open-source LLMs flourish, KD plays a crucial role in both compressing these models, and facilitating their self-improvement by employing themselves as teachers. This paper presents a comprehensive survey of KD’s role within the realm of LLM, highlighting its critical function in imparting advanced knowledge to smaller models and its utility in model compression and self-improvement. Our survey is meticulously structured around three foundational pillars: algorithm, skill, and verticalization – providing a comprehensive examination of KD mechanisms, the enhancement of specific cognitive abilities, and their practical implications across diverse fields. Crucially, the survey navigates the intricate interplay between data augmentation (DA) and KD, illustrating how DA emerges as a powerful paradigm within the KD framework to bolster LLMs’ performance. By leveraging DA to generate context-rich, skill-specific training data, KD transcends traditional boundaries, enabling open-source models to approximate the contextual adeptness, ethical alignment, and deep semantic insights characteristic of their proprietary counterparts. This work aims to provide an insightful guide for researchers and practitioners, offering a detailed overview of current methodologies in knowledge distillation and proposing future research directions. By bridging the gap between proprietary and open-source LLMs, this survey underscores the potential for more accessible, efficient, and powerful AI solutions. Most importantly, we firmly advocate for compliance with the legal terms that regulate the use of LLMs, ensuring ethical and lawful application of KD of LLMs. An associated Github repository is available at https://github.com/Tebmer/Awesome-Knowledge-Distillation-of-LLMs.
@@ 1 Introduction
Primarily, recognizing the disparities between proprietary and open-source LLMs, KD techniques have surged as a means to bridge the performance gap between these models Gou et al. (2021); Gupta and Agrawal (2022). Knowledge distillation, in this context, involves leveraging the more advanced capabilities of leading proprietary models like GPT-4 or Gemini as a guiding framework to enhance the competencies of open-source LLMs. This process is akin to transferring the ‘knowledge’ of a highly skilled teacher to a student, wherein the student (e.g., open-source LLM) learns to mimic the performance characteristics of the teacher (e.g., proprietary LLM). Compared to traditional knowledge distillation algorithms Gou et al. (2021), data augmentation (DA) Feng et al. (2021) has emerged as a prevalent paradigm to achieve knowledge distillation of LLMs, where a small seed of knowledge is used to prompt the LLM to generate more data with respect to a specific skill or domain Taori et al. (2023). Secondly, KD still retains its fundamental role in compressing LLMs, making them more efficient without significant loss in performance.  Gu et al. (2024); Agarwal et al. (2024). More recently, the strategy of employing open-source LLMs as teachers for their own self-improvement has emerged as a promising approach, enhancing their capabilities significantly Yuan et al. (2024a); Chen et al. (2024a). Figure 1 provides an illustration of these three key roles played by KD in the context of LLMs.
Figure 1: KD plays three key roles in LLMs: 1) Primarily enhancing capabilities, 2) offering traditional compression for efficiency, and 3) an emerging trend of self-improvement via self-generated knowledge.
@@
A key aspect of the knowledge distillation is the enhancement of skills such as advanced context following (e.g., in-context learning Huang et al. (2022a) and instruction following Taori et al. (2023)), improved alignment with user intents (e.g., human values/principles Cui et al. (2023a), and thinking patterns like chain-of-thought (CoT) Mukherjee et al. (2023)), and deeper language understanding (e.g., machine reasoning Hsieh et al. (2023), semantic understanding Ding et al. (2023a), and code generation Chaudhary (2023)). These skills are crucial for the wide array of applications that LLMs are expected to perform, ranging from casual conversations to complex problem-solving in specialized domains. For instance, in vertical domains like healthcare Wang et al. (2023a), law LAW (2023), or science Zhang et al. (2024), where accuracy and context-specific knowledge are paramount, knowledge distillation allows open-source models to significantly improve their performance by learning from the proprietary models that have been extensively trained and fine-tuned in these areas.
@@
The benefits of knowledge distillation in the era of LLMs are multifaceted and transformative Gu et al. (2024). Through a suite of distillation techniques, the gap between proprietary and open-source models is significantly narrowed Chiang et al. (2023); Xu et al. (2023a) and even filled Zhao et al. (2023a). This process not only streamlines computational requirements but also enhances the environmental sustainability of AI operations, as open-source models become more proficient with lesser computational overhead. Furthermore, knowledge distillation fosters a more accessible and equitable AI landscape, where smaller entities and individual researchers gain access to state-of-the-art capabilities, encouraging wider participation and diversity in AI advancements. This democratization of technology leads to more robust, versatile, and accessible AI solutions, catalyzing innovation and growth across various industries and research domains.
@@ 2.1 Comparing Traditional Recipe
Self-Align Sun et al. (2024b), WizardLM Xu et al. (2023a), WizardCoder Luo et al. (2023a),
WizardMath Luo et al. (2023b), AugGPT Dai et al. (2023a), TDG He et al. (2023b), leaf, text width=31em ] ] [ Curation [ UltraChat Ding et al. (2023b), Phi-1 Gunasekar et al. (2023), Phi-1.5 Li et al. (2023a),
@@
GKD Agarwal et al. (2024), QuantGPT Tao et al. (2022a), LLM-QAT Liu et al. (2023a), , leaf, text width=31em ] ] [ Feedback [ CAI Bai et al. (2022a), WizardMath Luo et al. (2023b), UltraFeedback Cui et al. (2023a),
@@
TED Liang et al. (2023a), GKD Agarwal et al. (2024),BabyLlamaTimiryasov and Tastet (2023), leaf, text width=31em ] ] [ Reinforcement Learning [ CAI Bai et al. (2022a), UltraFeedback Cui et al. (2023a), WizardMath Luo et al. (2023b),
@@
WizardMath Luo et al. (2023b), Llama-GPT4 Peng et al. (2023a), , leaf, text width=31em ] ] [ Multi-turn Dialogue [ Vicuna Chiang et al. (2023), Baize Xu et al. (2023b), UltraLLaMA Ding et al. (2023b),
CAMEL Li et al. (2023b), OpenChat Wang et al. (2023c), Zephyr Tunstall et al. (2023), , leaf, text width=31em ] ] [ RAG Capbility [ KARD Kang et al. (2023a), SAIL Luo et al. (2023c), Self-RAG Asai et al. (2023), , leaf, text width=31em ] ] ] [ Alignment [ Thinking Pattern [ Selfee Ye et al. (2023), Orca Mukherjee et al. (2023), Orca 2 Mitra et al. (2023),
@@
Zephy Tunstall et al. (2023), UltraFeedback Cui et al. (2023a), , leaf, text width=31em ] ] [ Value [ CAI Bai et al. (2022a), Align Honesty Yang et al. (2023a), SANDBOX Liu et al. (2023b),
@@
RankZephyr Pradeep et al. (2023b), ExaRanker Ferraretto et al. (2023), , leaf, text width=31em ] ] [ Recommendation [ NDR Mysore et al. (2023), InstrcutRec Zhang et al. (2023b), ONCE Liu et al. (2023c), , leaf, text width=31em ] ] [ Text Generation Evaluation [ PandaLM Wang et al. (2023b), Prometheus Kim et al. (2024), InstructScore Xu et al. (2023d),
@@
Phi-1 Gunasekar et al. (2023), PERsD Chen et al. (2023a), MFTCoder Liu et al. (2023d),
WaveCoder Yu et al. (2024), Code Clean Jain et al. (2023), , leaf, text width=31em ] ] ] [ Multi-Modality [ LLaVA Liu et al. (2023e), SVIT Zhao et al. (2023b), LVIS-Instruct4V Wang et al. (2023e), Shikra Chen et al. (2023c),
@@
2.2 Relation to Data Augmentation (DA)
2.3 Survey Scope
Skill Distillation. This facet examines the specific competencies and capabilities enhanced through KD. It encompasses detailed discussions on context following Taori et al. (2023); Luo et al. (2023c), with subtopics like instruction following and retrieval-augmented generation (RAG) Capability. In the realm of alignment Mitra et al. (2023); Tunstall et al. (2023), the survey investigates thinking patterns, persona/preference modeling, and value alignment. The ‘agent’ category delves into skills such as Tool Using and Planning. NLP task specialization Dai et al. (2023a); Jung et al. (2023); Chaudhary (2023) is scrutinized through lenses like natural language understanding (NLU), natural language generation (NLG), information retrieval, recommendation systems, text generation evaluation, and code generation. Finally, the survey addresses multi-modality Liu et al. (2023e); Zhao et al. (2023b), exploring how KD enhances LLMs’ ability to interpret and integrate multiple forms of input, enriching their utility and applicability across various contexts.
@@ 3.1.2 Expansion
While the labeling approach is simple and effective, it faces certain limitations. Primarily, it is constrained by the scale and variety of the input data. In real-world applications, especially those involving user conversations, there are also concerns regarding the privacy of the data involved. To address these limitations, various expansion methods have been proposed Wang et al. (2022a); Taori et al. (2023); Chaudhary (2023); Si et al. (2023); Ji et al. (2023a); Luo et al. (2023b, a); Wu et al. (2023c); Sun et al. (2024b); Xu et al. (2023a); Guo et al. (2023c); Rozière et al. (2023); West et al. (2022). These methods take the demonstrations as seed knowledge and aim to expand a large scale and various data by in-context learning.
@@
Expansion techniques have been widely utilized to extract extensive instruction-following knowledge from teacher LLMs. Wang et al. (2022a) first introduces an iterative bootstrapping method, Self-Instruct, to utilize LLMs to generate a wide array of instructions based on several demonstrations sampled from 175 manually-written instructions. The newly generated instructions are then added back to the initial pool, benefiting subsequent expansion iterations. Subsequently, Taori et al. (2023) applies this expansion method to a more powerful teacher LLM, text-davinci-003, to distill 52K high-quality data. To improve the diversity and coverage during expansion, Wu et al. (2023c) and Sun et al. (2024b) prompt the teacher LLM to generate instructions corresponding to some specific topics. Xu et al. (2023a) propose an Evol-Instruct method to expand the instructions from two dimensions: difficulty (e.g. rewriting the question to be more complex) and diversity (e.g. generating more long-tailed instructions). This Evol-Instruct method is domain-agnostic and has been used to expand the distillation of coding Luo et al. (2023a) and math Luo et al. (2023b). Additionally, expansion methods can significantly augment NLP task datasets with similar samples, thereby enhancing task performance. For instance, AugGPT Dai et al. (2023a) leverages a teacher LLM to rephrase each sentence in the training samples into multiple conceptually similar, but semantically varied, samples to improve classification performance. Similarly, TDG He et al. (2023b) proposes the Targeted Data Generation (TDG) framework, which automatically identifies challenging subgroups within data and generates new samples for these subgroups using LLMs through in-context learning.
@@ 3.1.3 Data Curation
The pursuit of high-quality and scalable data generation in knowledge distillation from LLMs has led to the emergence of the Data Curation approach. This method arises in response to the limitations observed in both the Labeling and Expansion approaches. These methods often yield data of variable quality and face constraints in quantity. In Labeling, the seed knowledge is sourced from task datasets, leading to potential noise and dirty data. Meanwhile, in Expansion, the input $x$ is derived from seed demonstrations, which can result in homogeneous data when generated in large quantities. To overcome these challenges, the Data Curation method curates high-quality or large-scale data by extensive meta-information as seed knowledge Ding et al. (2023b); Gunasekar et al. (2023); Li et al. (2023a); Mar (2023); Liu et al. (2023d); Wei et al. (2023); Yu et al. (2024); Ye et al. (2022); Gao et al. (2023a); Yang and Nicolai (2023).
@@
Different studies primarily vary in their source and method of leveraging meta-information. UltraChat Ding et al. (2023b) effectively demonstrates the process of curating both high-quality and diverse data by distilled knowledge. They collect extensive meta-information across three domains: Questions about the World, Creation and Generation, and Assistance on Existing Materials. For example, under Questions about the World, they explore 30 meta-topics like ”Technology” and ”Food and Drink.” the teacher LLMs then use this meta-information to distill a broad array of instructions and conversations, achieving a substantial scale of 1.5 million instances. UltraChat stands out with its lexical and topical diversity. The UltraLLaMA model, fine-tuned on this data, consistently surpasses other open-source models. Another notable series, phi Gunasekar et al. (2023); Li et al. (2023a); Mar (2023), focuses on distilling smaller, high-quality datasets akin to ”textbooks.” Phi-1Gunasekar et al. (2023) experiments with synthesizing ”textbook quality” data in the coding domain. Their approach involves distilling clear, self-contained, instructive, and balanced content from LLMs, guided by random topics or function names to enhance diversity. The distilled data is a synthesis of 1 billion tokens of Python textbooks, complete with natural language explanations and code snippets, as well as 180 million tokens of Python exercises with solutions. Remarkably, the phi-1 model, despite its smaller size, outperforms nearly all open-source models on coding benchmarks like HumanEval and MBPP while being 10 times smaller in model size and 100 times smaller in dataset size. MFTCoder Liu et al. (2023d) utilizes hundreds of Python knowledge points as meta-information to create a CodeExercise Dataset. In contrast, Magicoder Wei et al. (2023) and WaveCoder Yu et al. (2024) get raw code collections from open-source code datasets, using this as meta-information for generating instructional data. In the context of NLU tasks, certain studies Ye et al. (2022); Gao et al. (2023a); Wang et al. (2021a) explore the use of labels as meta-information to synthesize corresponding samples for data augmentation. Similarly, in information retrieval tasks, there are efforts to utilize documents as meta-information for generating potential queries, thereby constructing large-scale retrieval pairs Bonifacio et al. (2022); Meng et al. (2023).